OpenStudy (anonymous):

4 5/6 + w + 2 3/4=12 A. 5 5/12 B. 5 1/5 C. 4 7/12 D. 4 5/12

OpenStudy (anonymous):

@chichappenedalexis @inowalst @LeibyStrauss @godmod360 @appleduardo @TheGaMer007 @ziggy12 @quickstudent

OpenStudy (anonymous):

What do you think the answer is? Any guesses? I'll give you a hint, it's not B.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Oh, and let me clarify something. Is your equation 45/6 + w + 23/4 = 12 or, 4 and 5/6 + w + 2 and 3/4 = 12?

OpenStudy (anonymous):

Mmm, okay well some people are tagging me so I have to go, But i'm just assuming you meant 4 and 5/6 + w + 2 and 3/4 = 12. So, that would be D. 4 and 5/12. Medal?

OpenStudy (anonymous):

ok so what would be the answer for that was i correct

OpenStudy (anonymous):

You were super close! It's not C it's D.